OpenStudy (anonymous):

Find the area of a square if the perimeter is equal to 38 in.

hartnn (hartnn):

perimeter = 4* side find side from here. then area is just A= side^2.

OpenStudy (anonymous):

9.5?

hartnn (hartnn):

yes, side = 9.5 Area = 9.5^2 = ... ?

OpenStudy (anonymous):

90.25?

hartnn (hartnn):

thats correct area . :)

hartnn (hartnn):

write it as 90.25 sq. in

OpenStudy (anonymous):

ok!